1. Find all values of m so that the function y = e^mx is a solution of the given differential equation. (Enter your answers as a comma-separated list.)

y prime + 3y = 0

m =____

2. Determine whether Theorem 1.2.1 guarantees that the differential equation
y prime = square root y^2 -25 possesses a unique solution through the given point.
(1, 7)

A. Yes
B. No
1
Expert's answer
2020-06-18T20:21:47-0400

1 "y = e^{mx}" is solution of "\frac{dy}{dx} = 3y"

So differentiate y with respect to x,

"\frac{dy}{dx} = me^{mx}"


now, as per equation, "(m+3)e^{mx} = 0"

solution for this is "m=-3" only.



2 Given differential equation,


"\frac{dy}{dx} = \sqrt {y^2 - 25}"


Solving this equation,


"\int \frac{dy}{\sqrt {y^2-25}} = \int dx"



"ln|y+\sqrt {y^2-25}| = x + lnC" . . . . . . . . (i)


At point (1,7), value of constant C,


"ln|7+\sqrt{49-25}| = 1 + lnC"


"lnC = ln|7+\sqrt{24}| -1"



putting back value in equation (i)



"ln|y+\sqrt{y^2 - 25}| = x + ln|7+\sqrt{24}|-1"


This equation is unique solution for differential equation.